OpenStudy (anonymous):

find all real and complex zeros of p(x)=x^3-x^2-4x-6

OpenStudy (nurali):

p(x) = x³ - x² - 4x - 6 = (x - 3) (x² + 2x + 2) ........ will be 0 if either factor is 0 x - 3 = 0 .... ⇒ x = 3 x² + 2x + 2 = 0 (x + 1)² = -1 .... ⇒ x = -1 ± i Answer: x = 3, -1 ± i